OpenStudy (periodicninja):

Evaluate fourth root of 9 multiplied by square root of 9 over the fourth root of 9 to the power of 5.

OpenStudy (periodicninja):

@Jadeishere

OpenStudy (periodicninja):

@TheSmartOne

OpenStudy (periodicninja):

i dont know where to even start with this equation

OpenStudy (jadeishere):

\[\frac{ \sqrt[4]{9} \times \sqrt{9}}{ \sqrt[4]{9}^2 }\] Does it look like this?

OpenStudy (jadeishere):

Show your work. That does not help in any way :)

OpenStudy (periodicninja):

it's not even one of the answers

OpenStudy (jadeishere):

Okay then :) What we can do is rewrite the terms :) \[\sqrt[4]{9}\] can be written as \[9^{1/4}\] \[\sqrt{9} = 3\] \[\sqrt[4]{9}^5\] can be rewritten as \[9^{5/4}\]

OpenStudy (jadeishere):

Any questions so far, as to why it does that?

OpenStudy (periodicninja):

what happened to the fact that the two parts where being divided?

OpenStudy (jadeishere):

Oh, they still are, I'm just breaking down the terms and rewriting them

OpenStudy (periodicninja):

ok

OpenStudy (jadeishere):

But, do you understand why they're changing into the new form? without the radical

OpenStudy (periodicninja):

it looks like you're treating the square root sigh as if it was a = sign and moving the powers to the other side with division

OpenStudy (jadeishere):

Yes, pretty much :) So, when you're looking for the root of a number, whether it's a square root, cubed root, or any other root, it's to a fractional power. Now we have this; \[\frac{ 9^{1/4} \times3 }{ 9^{5/4} }\]

OpenStudy (periodicninja):

where did three come from?

OpenStudy (jadeishere):

the square root of 9 is 3. 3 multiplied by 3 is equal to 9 :)

OpenStudy (periodicninja):

ok

OpenStudy (periodicninja):

then what

OpenStudy (jadeishere):

okay, sorry, I'm helping multiple people right now >.<

OpenStudy (jadeishere):

Then, since the base of the exponents are the same we subtract (as done when dividing exponents) the exponents, always the bigger number by the larger so \[9^{5/4} - 9^{1/4} = 9^{4/4} = 9\] then we end up with 3/9, which equals 1/3. Obviously, something is incorrect, so what we can do is make the 3 back into the square root of 9. \[\sqrt{9} = 9^{1/2}\] now we have \[\frac{ 9^{1/2} }{ 9 }\]

OpenStudy (jadeishere):

transforming 9 into \[9^{1} \] then subtract \[9^{1/1} - 9^{1/2}\]

OpenStudy (jadeishere):

@agent0smith I think I did something wrong, please help!!

OpenStudy (robtobey2):

\[\frac{\sqrt[4]{9} \sqrt{9}}{9^{5/4}}=\frac{9^{3/4}}{9^{5/4}}=9^{\frac{3}{4}-\frac{5}{4}}=9^{-2/4}=\frac{1}{\sqrt{9}}=\frac{1}{3} \]

OpenStudy (agent0smith):

@Jadeishere \[\large \frac{ 9^{1/2} }{ 9^1 } = 9^{\frac{ 1 }{ 2} - 1}\]

OpenStudy (jadeishere):

ohhhh, Okay.